How many ways can be letters { 5.a. 4.b, 1.c} be arranged so that all letters of

same kind are in a single block


1
Expert's answer
2020-08-17T18:56:53-0400

Number of ways of selecting 3 different letters

"=3C_3 = 1"

Number of ways to select 1 similar and 2 different letter:

"=2C_1\times 2C_2 = 2"

Number of ways of selecting 1 similar + 1 more similar letter and 1 different letter:

"=2C_2\times 1C_1 = 1"

Number of ways to select 2 similar and 1 different letter:

"=2C_2\times 2C_1 = 2"

ways of selecting 3 similar letters:

"=" 1

total ways:  "=1+2+1+2+1 =7"